OpenStudy (anonymous):

f(x)=1-x^2 is this a function

hero (hero):

@christinashalee, have you tried graphing 1 - x^2 yet?

OpenStudy (anonymous):

no not yet

OpenStudy (anonymous):

Graph it and use the vertical line test for a function

OpenStudy (anonymous):

ok thanks

OpenStudy (mosaic):

You can graph the function mentally if you remember what the graph of x^2 looks like. y = x^2 is a parabola with vertex at the origin and opens upward. y = -x^2 will be a parabola with vertex at the origin that opens downward. y = 1-x^2 is the previous parabola that is shifted up by 1 unit: vertex at (0,1) and downward opening parabola. If you draw a vertical line at any x value will it intersect the parabola just once or more than once?

hero (hero):

If you're having trouble graphing it, you can use this program to help you: http://www.desmos.com